# MTVQA: Benchmarking Multilingual Text-Centric Visual Question Answering

### v2

@@ Abstract
Text-Centric Visual Question Answering (TEC-VQA) in its proper format not only facilitates human-machine interaction in text-centric visual environments but also serves as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Nonetheless, most existing TEC-VQA benchmarks have focused on high-resource languages like English and Chinese. Despite pioneering works to expand multilingual QA pairs in non-text-centric VQA datasets through translation engines, the translation-based protocol encounters a substantial “visual-textual misalignment” problem when applied to TEC-VQA. Specifically, it prioritizes the text in question-answer pairs while disregarding the visual text present in images. Moreover, it fails to address complexities related to nuanced meaning, contextual distortion, language bias, and question-type diversity. In this work, we tackle multilingual TEC-VQA by introducing MTVQA, the first benchmark featuring high-quality human expert annotations across 9 diverse languages, consisting of 6,778 question-answer pairs across 2,116 images. Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA dataset, it is evident that there is still a large room for performance improvement, underscoring the value of MTVQA. Additionally, we supply multilingual training data within the MTVQA dataset, demonstrating that straightforward fine-tuning with this data can substantially enhance multilingual TEC-VQA performance. We aspire that MTVQA will offer the research community fresh insights and stimulate further exploration in multilingual visual text comprehension. The project homepage is at https://bytedance.github.io/MTVQA/.
footnotetext: Equal contribution. 🖂 Corresponding authors.
@@ 1 Introduction
In the era of burgeoning AI, especially in MLLMs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], Text-Centric Visual Question Answering (TEC-VQA[19, 20, 11, 10, 14, 13, 12] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Compared with general VQA [19, 21, 22, 20, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], TEC-VQA places greater emphasis on answering questions that require understanding textual information within images. It provides a streamlined avenue for individuals without specialized expertise to articulate their requirements and access applications in text-centric visual environments. However, the majority of advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English [19, 20, 21, 24], Chinese [35, 32], Japanese [36, 37] and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
To tackle the problem of language diversity, several seminal studies [38, 39, 40] in the general VQA field, leverage off-the-shelf translation engines to expand existing question-answer pairs from high-resource languages to their multilingual counterparts including low-resource ones. However, when applied to TEC-VQA, this translation-based approach may fall prey to the “visual-textual misalignment” problem as only the text in question-answer pairs can be processed, while the visual text present in the images is overlooked. Although the visual text can be obtained by an OCR (Optical Character Recognition) engine and then translated into English, this indirect process could introduce issues such as nuanced meaning, contextual distortion, language bias, and question type diversity. These factors further render the transferability of the translation protocol infeasible for TEC-VQA. Taking the second case in Fig. 1 as an example, if either the visual text or the question in Russian is recognized or translated problematically, then the question would never be answered. The status quo begs for a question: “Can we address the visual-textual misalignment problem directly from the source language per se for multilingual TEC-VQA and what we stand in the MLLM era?
Figure 1: Multilingual text-centric VQA visualization selected from four languages. From left to right: Arabic (AR), Russian (RU), Thai (TH), Vietnamese (VI). The corresponding translations in English are in brackets. More examples see Figure 7.
@@
We further investigate recent representative MLLMs in Fig. 3, including GPT-4o, GPT-4V, Gemini, QwenVL, etc., and their performance on our proposed MTVQA. As vividly revealed in Fig. 3 (a), GPT-4o is the top performer, followed by Claude3 Opus and Gemini Ultra, with closed-source MLLMs generally leading. General MLLMs like QwenVL Max and QwenVL Plus show mid-tier performance, while text-centric MLLMs lag behind. Fig. 3(b) shows closed-source MLLMs consistently outperforming others, especially in French (FR) and German (DE), while languages like Arabic (AR) and Thai (TH) witness lower scores across all categories. The results unequivocally demonstrate that opportunities for improvement persist within existing MLLMs when applied in multilingual text-rich scenarios. In summary, the main contributions of this paper can be categorized into three points:
  • We coin the MTVQA dataset, which to the best of our knowledge is the first multilingual TEC-VQA benchmark providing human expert annotations to solve “visual-textual misalignment” problem in multilingual text-centric scenarios.
  • We benchmark the state-of-the-art MLLMs on our new dataset and show that there are still opportunities for improvement on even the most advanced MLLMs when it comes to multilingual text-rich scenarios.
  • We establish a set of new multilingual TEC-VQA baselines for both closed-source and general-purpose, text-centric MLLMs.
@@ 2.1 MLLMs for Text-centric VQA
Recent advancements in MLLMs [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models. Notably, the high generalizability of MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [11, 10, 14, 13, 12]. For example, LLaVAR [41], UniDoc [11], which extend LLaVA [31] into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images. Furthermore, DocPedia [10] operates visual input in the frequency domain rather than in space, which enables higher input resolution without increasing the input sequence. Lately, mPLUG-DocOwl [42], Qwen-VL [7], and TextMonkey [13] leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities. Despite the promising results achieved by existing MLLMs in text-centric VQA tasks, their focus on high-resource languages such as English and Chinese has posed challenges in achieving reasonable performance for low-resource languages. This is primarily due to the lack of data or benchmarks for these low-resource languages.
@@ 2.2 Multilingual Text-centric VQA Benchmarks
There are some recent efforts toward extending VQA tasks to a wider range of languages [43, 39, 44, 40, 45, 38] by providing a multilingual VQA datasets. For example,  Gao et al. [32] created a free-form bilingual VQA dataset (FM-IQA) containing over 150,000 images and 310,000 freestyle Chinese question-answer pairs and English translations. Raj Khan et al. [38] developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages. Of more relevance are the works xGQA (8 languages) [39] and MaXM (7 languages) [40], which apply translation-based protocols to expand VQA data beyond English. However, the translation-based multilingual VQA datasets inherently face issues, such as the “visual-textual misalignment” problem, where only the textual information in question-answer pairs is considered, while the visual text in images is overlooked. Additionally, the nuanced meaning and context are often distorted; language bias is introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by [40]. Moreover, none of the previous multilingual datasets focus on text-centric scenarios where multilingual text frequently occurs.
Our benchmark, MTVQA, distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations. It covers 9 languages, thereby facilitating the training and evaluation of multilingual models in diverse linguistic contexts. Additionally, our dataset can gauge the VQA system’s ability for not only high-resource languages but also those that are typically underrepresented in current datasets [19, 20, 21, 24, 32].
@@ 3 MTVQA Benchmark
The MTVQA benchmark is meticulously established to evaluate the multilingual text comprehension performance of Multimodal Large Language Models (MLLMs). MTVQA covers nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). The data construction process involves a comprehensive collection of text-rich images and a two-round human expert annotation process.
Figure 4: A brief diagram of the annotation process.
@@ 3.1 Data Collection
The raw data collection is oriented towards gathering text-rich images from various natural and document scenarios, ensuring diversity and quality. This includes images from publicly available datasets(e.g., ICDAR MLT19 [46]) and those sourced from the internet, such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig. 2. A total of 1,220 images from document scenarios and 876 images from natural scenarios are collected for the test set of the MTVQA benchmark. To ensure the visual-textual alignment, for text-rich images lacking text and language annotations, we subject them to a standardized data cleaning process, which includes multilingual text recognition and language classification. Afterwards, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
@@ 3.2 Human Expert Annotation
To obtain informative and accurate text-related QA pairs for images grouped by specific languages, a specialized group of annotators with expertise in the local regions of each language is recruited. These annotators are native speakers, guaranteeing a profound understanding and skillfulness in the linguistic subtleties and cultural contexts required for precise annotations. Given the subjective nature of understanding text within images, the annotation team is divided into two independent groups. One group is tasked with generating questions and providing answers based on the images, while the other group is responsible for evaluating and correcting the QA pairs. This division, known as the raise-then-correct paradigm, ensures a thorough and trustworthy evaluation of the text-rich image comprehension process. Moreover, the annotation results for each language underwent a 10% sampling inspection by a quality inspector to ensure adherence to standards. QA pairs that are not meeting the criteria are returned for re-annotation. Before beginning the formal annotation task, all annotators have received training and examples of annotations. The two-round annotation process is briefly illustrated in Fig. 4, further details of which are elaborated in the subsequent sections.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Word clouds showcasing top answers in various languages, tokenized via NLTK with removing stop words, punctuation, and digits.
@@
Second Round Evaluation and Correction. To reduce the effect of human subjective cognitive bias on our MTVQA benchmark and get high-quality question-answer pairs, we assign two annotators for each language to carry out the annotation evaluation and correction process. These annotators follow a specific evaluation and correction protocol to ensure consistency and accuracy based on the first round results: (1) Assess the relevance of the question to the text in the image. Irrelevant QA pairs are discarded; (2) Verify the correctness of the answer and make necessary modifications; (3) Check for redundancy in the answer. If the answer repeats the question’s information, the repeated content is removed for conciseness. (4) Ethical Assessment. Review the image content or the question-answer pair for any unethical content, including but not limited to politics, personal privacy issues, etc. Such content is removed to uphold ethical standards in the dataset.
@@ 4.1 Baseline Models
To comprehensively assess MLLMs’ multilingual perception and comprehension capabilities, We select state-of-the-art MLLMs from three categories: open-source general MLLMs, open-source text-centric MLLMs, and closed-source MLLMs. Each category contains the following models: (1) General MLLMs: InternVL-V1.5 [47], InternLM-Xcomposer2-4KHD [16], Mini-Gemini-HD-34B [17], Llava-Next-34B [18], DeepSeek-VL [8], YI-VL-34B [9]; (2) Text-centric MLLMs: TextSquare [14], TextMonkey [13], mPLUG-DocOwl 1.5 [12], MiniCPM-V 2.0 [48]; (3) Closed-source MLLMs: GPT-4o [1], GPT-4V [2], Gemini Ultra [4], QwenVL Max [7], QwenVL Plus [7], Claude3 Opus [5], Claude3 Sonnet [5], and GLM-4V [49]. For the closed-source MLLMs, we use the chat version through official APIs, while for the open-source MLLMs, we utilize the instruct versions available on the HuggingFace Model Hub. The open-source MLLMs’ model size varies from 7b to 34b.
4.2 Implementation Details
Table 1: Performance of the leading closed- and open-source MLLMs on the MTVQA benchmark. The best results of each language are bolded. The best results for open-source models are underlined. “Xcomposer-SFT” denotes instruction tuning to Xcomposer2-4KHD with MTVQA’s training set.
@@ 4.2 Implementation Details
4.3 Evaluation Results
Evaluation metric. To accurately assess whether the visual text that occurs in the answer is correct, we adopt Accuracy as the metric. The Accuracy metric measures the percentage of questions for which the predicted answer matches exactly with any of the target answers for the question. The accuracy metric awards a zero score even when the prediction differs slightly from the target answer.
[…]

### v3

@@ Abstract
Text-Centric Visual Question Answering (TEC-VQA) in its proper format not only facilitates human-machine interaction in text-centric visual environments but also serves as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Nonetheless, most existing TEC-VQA benchmarks focus on high-resource languages like English and Chinese. Despite pioneering works expanding multilingual QA pairs in non-text-centric VQA datasets through translation engines, the translation-based protocol encounters a substantial “visual-textual misalignment” problem when applied to TEC-VQA. Specifically, it prioritizes the text in question-answer pairs while disregarding the visual text present in images. Moreover, it fails to address complexities related to nuanced meaning, contextual distortion, language bias, and question-type diversity. In this work, we tackle multilingual TEC-VQA by introducing MTVQA, the first benchmark featuring high-quality human expert annotations across 9 diverse languages, consisting of 6,778 question-answer pairs across 2,116 images. Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including Qwen2-VL, GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA benchmark, it is evident that there is still a large room for performance improvement (Qwen2-VL scoring 30.9 versus 79.7 for human performance), underscoring the value of MTVQA. Additionally, we supply multilingual training data within the MTVQA dataset, demonstrating that straightforward fine-tuning with this data can substantially enhance multilingual TEC-VQA performance. We aspire that MTVQA will offer the research community fresh insights and stimulate further exploration in multilingual visual text comprehension.
@@ 1 Introduction
In the era of burgeoning AI, especially in MLLMs [34, 1, 47, 44, 3, 39, 5, 27, 49, 12, 13, 18, 26, 43, 9, 11, 22, 24], Text-Centric Visual Question Answering (TEC-VQA[6, 42, 13, 12, 43, 26, 18] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Compared with general VQA [6, 30, 36, 42, 32, 31, 29, 52, 20, 4, 28, 40, 25, 15, 14, 23], TEC-VQA places greater emphasis on answering questions that require understanding visual textual information within images. It enables individuals without specialized expertise to access applications in text-centric visual environments. However, most advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English [6, 42, 30, 31], Chinese [37, 15], Japanese [41, 33] and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
To tackle the problem of language diversity, several seminal studies [38, 35, 7] in the general VQA field simply leverage translation engines to expand question-answer pairs from high-resource to low-resource languages. However, this translation-based approach is not feasible for TEC-VQA, as it merely processes text in question-answer pairs, neglecting the critical visual text needed for answering based on images. Although the visual text can be recognized by an OCR (Optical Character Recognition) engine and then translated into the target language, this indirect process could lead to a “visual-textual misalignment” problem, due to nuanced meaning, contextual distortion, language bias, and question type diversity. Taking the second case in Fig. 1 as an example, if either the visual text or the question in Russian is recognized or translated problematically, then the question would never be answered correctly. The status quo begs for a question: “Can we directly leverage visual text in source language per se for multilingual TEC-VQA and what we stand in the MLLM era?
Figure 1: Multilingual text-centric VQA visualization selected from four languages. From left to right: Arabic (AR), Russian (RU), Thai (TH), Vietnamese (VI). The corresponding translations in English are in brackets.
@@
We further investigate recent representative MLLMs in Fig. 3, including GPT-4o, GPT-4V, Gemini, Qwen2-VL, etc., and their performance on MTVQA. As vividly revealed in Fig. 3 (a), the general MLLM Qwen2-VL is the top performer, followed by closed-source GPT-4o and Claude3 Opus. Other general MLLMs like QwenVL Max and QwenVL Plus show mid-tier performance, while text-centric MLLMs lag behind. Fig. 3(b) shows closed-source MLLMs consistently outperforming others, especially in French (FR) and German (DE), while languages like Arabic (AR) and Thai (TH) witness lower scores across all categories. The results unequivocally demonstrate that opportunities for improvement persist within existing MLLMs when applied in multilingual text-rich scenarios. In summary, the main contributions of this paper can be categorized into four points:
  • We coin the MTVQA dataset, which, to the best of our knowledge, is the first multilingual TEC-VQA benchmark providing human expert annotations to solve the “visual-textual misalignment” problem in multilingual text-centric scenarios.
  • We benchmark various MLLMs and human performance on MTVQA and show that there are still great opportunities for improvement on even the most advanced MLLMs in multilingual text-rich scenarios.
  • We establish a set of new multilingual TEC-VQA baselines for closed-source and general-purpose, text-centric MLLMs.
  • MTVQA has been serving as a standard benchmark for evaluating multilingual visual text comprehension capabilities by some recently released state-of-the-art MLLMs such as Qwen2-VL [46] and InternVL 2 [11].
@@ 2 Related Work
Table 1: Comprehensive comparison on the benchmarks related to MTVQA. GPT4V denotes GPT-4V performance.
Benchmark Scene Manual QA QA Language Visual Text Language GPT4V
OCRBench Multiple Text-rich English English 64.5%
TextVQA Scene Text English English 78.0%
DocVQA Document English English 88.4%
EST-VQA Scene Text English, Chinese English, Chinese 72.3%
xGQA General 7 Languages - 67.7%
MaXM General 7 Languages - 62.8%
MTVQA Multiple Text-rich 9 Languages 10 languages 22.0%
@@ 2.1 MLLMs for Text-centric VQA
Recent advancements in MLLMs [1, 47, 44, 3, 39, 5, 27, 49, 12, 13, 18, 26, 43, 9, 11, 22, 24, 51] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models. Notably, the high generalizability of MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [13, 12, 43, 26, 18]. For example, LLaVAR [50], UniDoc [13], which extend LLaVA [25] into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images. Furthermore, DocPedia [12] operates visual input in the frequency domain rather than in space, which enables higher input resolution without increasing the input sequence. Lately, mPLUG-DocOwl [48], Qwen-VL [5], and TextMonkey [26] leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities. Despite the promising results achieved by existing MLLMs in text-centric VQA tasks, their focus on high-resource languages such as English and Chinese has posed challenges in achieving reasonable performance for low-resource languages. This is primarily due to the lack of data or benchmarks for these low-resource languages.
@@ 2.2 Multilingual Text-centric VQA Benchmarks
There are some recent efforts toward extending VQA tasks to a broader range of languages [17, 35, 45, 7, 21, 38] by providing a multilingual VQA datasets. For example,  Gao et al. [15] created a free-form bilingual VQA dataset (FM-IQA) containing over 150,000 images and 310,000 freestyle Chinese question-answer pairs and English translations. Raj Khan et al. [38] developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages. Of more relevance are the works xGQA (7 languages) [35] and MaXM (7 languages) [7], which apply translation-based protocols to expand VQA data beyond English. However, the translation-based multilingual VQA datasets inherently face issues, such as the “visual-textual misalignment” problem, where only the textual information in question-answer pairs is considered, while the visual text in images is overlooked. Additionally, the nuanced meaning and context are often distorted; language bias is introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by [7]. Moreover, none of the previous multilingual datasets focus on text-centric scenarios where multilingual text frequently occurs.
Our benchmark, MTVQA, distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations. It covers 9 languages, facilitating the training and evaluation of multilingual models in diverse linguistic contexts. Additionally, our dataset can gauge the VQA system’s ability for not only high-resource languages but also those that are typically underrepresented in current datasets [6, 42, 30, 31, 15].
@@ 3 MTVQA Benchmark
The MTVQA benchmark is meticulously established to evaluate the multilingual text comprehension performance of Multimodal Large Language Models (MLLMs). MTVQA covers nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). The data construction involves a comprehensive collection of text-rich images and a two-round human expert annotation process. In terms of data volume, an initial dataset of 8,980 images is compiled. Following a data cleaning phase, 8,895 images are retained for annotation. The first round of annotation yielded 8,895 images, corresponding to 28,906 question-answer (QA) pairs. After the second round of annotation and final quality control measures, the final dataset comprises 8,794 images and 28,607 QA pairs. The benchmark construction cost is divided into two parts: image acquisition and annotation. Image acquisition costs about two months and 30,000 dollars. VQA annotation costs about three months and 60,000 dollars. Hourly wages vary from country to country, with senior language specialists costing between 20 and 40 per hour and local annotators costing between 8 and 20 per hour. The average time spent by each labeler was roughly 60 days.
Figure 4: A brief pipeline of the annotation process.
@@ 3.1 Data Collection
The raw data collection aims to gather text-rich images from various scene text and document scenarios, ensuring diversity and quality. This includes images from publicly available datasets (e.g., ICDAR MLT19 [16]) and those sourced from the internet (i.e., Laion-OCR, which is filtered from Common Crawl [10]), such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig. 2. The image collection process consists of two steps: extracting multilingual text using a multilingual OCR engine (Bytedance multilingual OCR API [^1] https://cv-api.bytedance.com/api/ocr-process/v1/multi_language_v2) and selecting by language types and the amount of text contained in the image. Approximately 30% of the overall image data is obtained from public datasets, with 20% sourced from the web and 50% from manual collection, respectively. Manual collection means that the images are taken in different scenarios in the countries and areas where the corresponding language is spoken, which means that the images are native and high quality. A total of 1,220 images from document scenarios and 876 images from natural scenarios are collected for the test set of the MTVQA benchmark. To ensure data content meets regulatory requirements, we subject them to a standardized data cleaning process. The image-cleaning processing pipeline involves a preliminary round of algorithm-driven review to identify and filter out unusable images. It includes detecting and removing images with politically sensitive, pornographic, violent, or other undesirable features. Subsequently, the multilingual OCR tool (Bytedance multilingual OCR API) is employed to extract the textual content from the remaining images. Images devoid of textual information are discarded, and the surviving images are categorized based on their language. Afterward, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
@@ 3.2 Human Expert Annotation
To obtain informative and accurate text-related QA pairs for images grouped by specific languages, a specialized group of annotators with expertise in the local regions of each language is recruited. The annotators must be native speakers of the corresponding language and have actively utilized it for a minimum duration of 10 years. Additionally, they must possess at least a university-level degree or higher academic qualification, guaranteeing a profound understanding and skillfulness in the linguistic subtleties and cultural contexts required for precise annotations. Given the subjective nature of understanding text within images, the annotation team is divided into two independent groups. One group is tasked with generating questions and providing answers based on the images, while the other group is responsible for evaluating and correcting the QA pairs. This division, known as the raise-then-correct paradigm, ensures a thorough and trustworthy evaluation of the text-rich image comprehension process. Moreover, the annotation results for each language underwent a 10% sampling inspection by a quality inspector to ensure adherence to standards. QA pairs that do not meet the criteria are returned for re-annotation. Prior to the formal annotation process, all annotators will be provided with a detailed explanation of the annotation guidelines, including a dedicated question-and-answer session to clarify any ambiguities and uncertainties. A pilot annotation task will be conducted on a limited dataset to ensure a shared understanding of the annotation rules. The two-round annotation process is briefly illustrated in Fig. 4, further details of which are elaborated in the subsequent sections.
@@
Second Round Evaluation and Correction. To reduce the effect of human subjective cognitive bias on our MTVQA benchmark and get high-quality question-answer pairs, we assign two annotators for each language to carry out the annotation evaluation and correction process independently. These annotators follow a specific evaluation and correction protocol to ensure consistency and accuracy based on the first round results: (1) Assess the question’s relevance to the text in the image. Irrelevant QA pairs are discarded; (2) Verify the correctness of the answer and make necessary modifications; (3) Check for redundancy in the answer. If the answer repeats the question’s information, the repeated content is removed for conciseness. (4) Ethical Assessment. Review the image content or the question-answer pair for any unethical content, including but not limited to politics, personal privacy issues, etc. Such content is removed to uphold ethical standards in the dataset.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ 4.1 Baseline Models
To comprehensively assess MLLMs’ multilingual perception and comprehension capabilities, We select state-of-the-art MLLMs from three categories: open-source general MLLMs, open-source text-centric MLLMs, and closed-source MLLMs. Each category contains the following models: (1) General MLLMs: Qwen2-VL  [46], InternVL-V1.5 [8], InternLM-Xcomposer2-4KHD [11], Mini-Gemini-HD-34B [22], Llava-Next-34B [24], DeepSeek-VL [27], YI-VL-34B [49]; (2) Text-centric MLLMs: TextSquare [43], TextMonkey [26], mPLUG-DocOwl 1.5 [18], MiniCPM-V 2.0 [19]; (3) Closed-source MLLMs: GPT-4o [34], GPT-4V [1], Gemini Ultra [44], QwenVL Max [5], QwenVL Plus [5], Claude3 Opus [3], Claude3 Sonnet [3], and GLM-4V [2]. For the closed-source MLLMs, we use the chat version through official APIs, while for the open-source MLLMs, we utilize the instruct versions available on the HuggingFace Model Hub. The open-source MLLMs’ model size varies from 7b to 76b.
models GPT-4V GPT-4o Gemini Ultra Claude3 Opus Qwen2-VL 92B InternVL2 76B
ANLS 29.6 36.9 31.7 35.4 38.7 29.1
Table 2: Performance of the leading closed- and open-source MLLMs on the MTVQA benchmark. The best results of each language are bolded. The second best results are underlined. “Xcomposer-SFT” denotes instruction tuning to Xcomposer2-4KHD with MTVQA’s training set.
@@
4.2 Implementation Details